OpenStudy (hugsnotughs):

I need some help please! What is the rate of change from x = 0 to x = pi over 2?

OpenStudy (hugsnotughs):

http://prntscr.com/6oewk7

OpenStudy (anonymous):

rate of change it has gone from -4 to 0 in pi/2= -4/(pi/2) = -pi/8